OpenStudy (anonymous):

A camping supply store carries tents usually priced from 68$ to 199$. the store is having a sale. what is the range of possible prices you could pay for a tent. ( all tents on sale 10% to 25% off )

OpenStudy (anonymous):

we could begin by calculating the cheaper tent with the greater discount

OpenStudy (anonymous):

I mean 68 - (25/100)68 =

OpenStudy (anonymous):

then the more expensive tent with the more little discount

OpenStudy (anonymous):

199 - (10/100) 199 =

OpenStudy (anonymous):

first 68 - 17 = 51 second 199 - 19,9 = 179,1

OpenStudy (anonymous):

RANGE = 51$ to 179,1$

OpenStudy (anonymous):

well It was a pleasure to help you Saludos desde Perú